Solve the given initial-value problem. --- x xo) -(-) XO
Answer #1

kinen that X= AX where A- 2 4 Y 6. Now, the eigenvalues of A are 2-1 4 co 0 => def (A-nI₂] =0 -1 6-1 -> (2-u6-47 +420, 7 12-At X2 e re CH-41 741) re 44 (A-41) 2 e e 20 44 se [ IT (A-4A)* (A-45) ? 2 1! 21 +---] 48 = e .-27 4t -t at ze [16) t -27 -2t- (1)-(****) =5 (2=-1-24 = -1-102-11 → ht x 4t e -11 5 = ) 1-27 -t e is the required son CS Scanned with CamScanner